OpenStudy (anonymous):

3RD TIME POSTING!!!!!! Determine whether y varies directly with x. if so, find the constant variation of k.? x l y 4 l 6.4 7 l 11.2 10 l 16 13 l 20.8

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@ashley_f97

OpenStudy (anonymous):

@kropot72

OpenStudy (oaktree):

I'm sorry you had to wait so long for an answer! The way to solve this is by dividing each y by its corresponding x and seeing if the numbers you get are the same every time. Since\[\frac{ 6.4 }{ 4 } = \frac{ 11.2 }{ 7 } = \frac{ 16}{ 10 } = \frac{ 20.8 }{ 13 } = 1.6\]We know that there is a direct variance. The equation is then\[y = 1.6x\]Makes sense?

OpenStudy (anonymous):

you deserve 5 medals for this thank you very much

OpenStudy (oaktree):

No problem! Just fan me or something and that way you'll be able to grab me on any other questions you have.